Assuming a Ingersoll-Rand bond has a coupon rate of 6 %, pays semi-annual interest payments, and has twelve years to maturity. If investors are willing to accept a 8% rate of return on bonds of similar quality, what is the value of the Ingersoll-Rand? Enter answer in positive up to two decimal points. (X,XXX.XX)
Answer #1

M = $1000, n = 12 * 2 = 24 semi-annual periods, C = 6% * $1000/2 = $30 semi-annually, i = 8%/2 = 4% semi-annually

P = 457.41 + 390.12

P = $847.53